# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v2

@@ E.5 Limitations of Outcome Reward Model
The performance is described in Liu et al. (2024b), with results presented in Table˜8. Although this model achieves an average score of 64.79—representing an improvement of 3.23 points over the Chain-of-Thought (CoT) method—it remains significantly inferior to the performance attained using a self-rewarding mechanism, which reaches 77.18. These results highlight the limitations of current outcome-based reward models in scenarios lacking ground-truth data, underscoring the need for the development of more robust and effective outcome reward models.
@@
E.6 Performance vs Fraction Ratio
Figure˜5 illustrates the performance trends of various instruction-tuned language models across different fraction ratios, which range from 0.1 to 0.8. Here, we also experimented with another model: Mistral-7B-Instruct (Jiang et al., 2023). Performance is reported as a percentage and plotted on the y-axis, while the fraction ratio is shown on the x-axis.
@@ F.1 Prompt Designation
The prompts are listed in Tables˜21, 19, 20 and 18.
F.2 Backbone
We list all model backbone as follows:
F.3 Baselines
We describe all baselines in experiments as follows:
  • CoH Liu et al. (2023a): A framework which systematically aligns language models with human feedback through a structured, iterative process, enhancing their performance in complex reasoning tasks.
  • SPIN Chen et al. (2024a): A self-play fine-tuning methodology that significantly enhances the performance of weaker language models, transforming them into robust and highly capable systems competitive with stronger counterparts.
@@ F.3 Baselines
F.4 GSM8K
We provide details for GSM8K experiments as follows:
Dataset.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜12 and Table˜13, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The detailed specifications of these four evaluation prompts are provided in Tables˜18, 21, 19 and 20.The prompt structure is consistently applied across all backbones and both prompts. Notably, for Llama-3.1-8B-Instruct, Qwen2.5-1.5B-Instruct and Qwen2-7B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weights of 3, 2, and 2 assigned to format-based criteria, respectively.
Scaling.
We offer more scaling figures in Figure˜6.
(a) Qwen2.5-1.5B-Instruct
(b) Qwen2.5-7B-Instruct
(c) Qwen2.5-14B-Instruct
(d) LlaMA3.1-8B-Instruct
(e) Qwen2-7B-Instruct
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
LatentSeek (Self) Qwen2-7B-Instruct 1024 2 1 L40 0.03 Adam 0.2 bf16 10
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
F.5 MATH-500
We provide details for MATH-500 experiments as follows:
Dataset.
The MATH-500 dataset, a curated subset of the MATH benchmark, serves as a robust resource for evaluating the mathematical reasoning capabilities of machine learning models. Sourced from the HuggingFace repository at https://huggingface.co/datasets/HuggingFaceH4/MATH-500 , it comprises 500 problems derived from the original MATH benchmark developed by OpenAI (Lightman et al., 2023). Encompassing a diverse array of mathematical topics and varying difficulty levels, MATH-500 provides a comprehensive and challenging testbed for assessing model performance in mathematical problem-solving.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜14 and Table˜15, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts. Notably, for Qwen2-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using prompt 1, the weight ratios are adjusted to 1:1:1:2. Additionally, for Qwen2-7B-Instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weight of 2 assigned to format-based criteria.
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
F.6 AIME2024
We provide details for AIME2024 experiments as follows:
Dataset.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜16 and Table˜17, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts.
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
@@ G.1 Forward Pass FLOPs Estimation
The total FLOPs for a single forward pass is estimated by summing the operational costs of its constituent components, as detailed in Table˜22. The total FLOPs for one forward pass is calculated as follows:
@@ G.3 Total FLOPs for LatentSeek
LatentSeek operates exclusively at test time on 1319 instances. Each iteration involoves two forward passes (one for generation, one for reward evaluation) and one partial backward pass through only the LM head. From Table˜4, the average number of iterations on GSM8K is 1.27.
@@ G.4 Efficiency Threshold Analysis
Appendix H Qualitative Analysis and Case Studies
@@
To gain a deeper understanding of model-generated responses, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021), using the Llama3.1-8B-Instruct model. These distributions are visualized in the word cloud shown in Figure˜7. The appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and unrelated proper names like “Mark” and “John”—suggests that the reasoning processes employed by language models can diverge significantly from human reasoning, despite generating superficially plausible outputs. Notably, consistent patterns emerge: the first generated tokens are often prepositions (e.g., “let”), the second tokens are primarily verbs (e.g., “find,” “solve”), and the third tokens are predominantly proper names, These findings suggest that optimal reasoning paths for language models may differ from human strategies and could be more effectively modeled in latent space.
We present several representative cases in Tables˜23, 24, 25, 26, 27, 28 and 29 to demonstrate the divergence between machine reasoning patterns and human cognition processes. These cases reveal that while our method (OURS) generates linguistically anomalous expressions such as "total downloads of downloads", "let need calculate to calculate" and "let’ll more understand it", it still achieves correct solutions. This suggests that even if the decoded tokens from our updated embeddings lack syntactic and semantic coherence, the logic and answer of the final generated output remain accurate. Notably, although the baseline model produces erroneous answers that appear fundamentally distinct from ours (OURS) in textual form, their latent representations exhibit structural proximity - through minimal adjustments to embeddings in the latent space, the original output (ORIGINAL MODEL) can be transfromed into correct solutions (OURS). These empirical evidences yield two critical implications. First, it suggests that the reasoning paths employed by machines are fundamentally distinct from those of humans. Second, while model errors may appear substantially deviant from correct answers in textual outputs, they may maintain significant proximity to valid solutions in latent space.
@@
Appendix I Computational Resources
@@
Appendix J The Use of Large Language Models (LLMs)

### v3

@@ E.5 Limitations of Outcome Reward Model
The performance is described in Liu et al. (2024a), with results presented in Table˜12. Although this model achieves an average score of 64.79—representing an improvement of 3.23 points over the Chain-of-Thought (CoT) method—it remains significantly inferior to the performance attained using a self-rewarding mechanism, which reaches 77.18. These results highlight the limitations of current outcome-based reward models in scenarios lacking ground-truth data, underscoring the need for the development of more robust and effective outcome reward models.
@@
E.6 Trends of Reward
(a) Qwen2.5-1.5B-Instruct
(b) Qwen2.5-7B-Instruct
(c) Qwen2.5-14B-Instruct
(d) LlaMA3.1-8B-Instruct
(e) Qwen2-7B-Instruct
Figure 6: Trends of reward on GSM8K dataset. Generation uses the prompt 2 (the Json prompt).
We plot the reward trends in Figure˜6, which empirically illustrates how the update dynamics guide the model from a low-reward region toward a higher-reward region. The reward curves are obtained by running LatentSeek on the GSM8K dataset, with generation constrained by a JSON-formatted prompt.
E.7 Comparison of Optimizing Initial-Stage Latent Tokens and Optimizing Middle-Stage Latent Tokens
Table 13: Comparison with Optimizing the Middle-Stage Tokens on GSM8K. The best preformances are highlighted in bold.
Model Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8K: Prompt 1
CoT 68.01 68.08 88.86 92.03 50.19 73.43
LatentSeek (middle) 74.68 69.07 89.31 92.49 65.21 78.15
LatentSeek 84.38 70.89 90.14 92.49 78.54 83.29
GSM8K: Prompt 2
CoT 65.20 15.31 66.41 91.81 69.07 61.56
LatentSeek (middle) 70.73 22.44 72.93 92.49 77.26 67.17
LatentSeek 80.21 44.20 85.06 92.72 83.70 77.18
One may wonder how performance changes when we instead optimize middle-stage tokens. Formally, this corresponds to the factorization
$\pi(\mathbf{x}\mid\mathbf{z},\mathbf{c})=\prod_{t=1}^{N_{1}}\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c})\prod_{t=N_{1}+1}^{N}\pi_{\text{LM-head}}(x_{t}\mid z_{t})\prod_{t=N+1}^{T}\pi(x_{t}\mid\mathbf{x}_{<t},\mathbf{c}).$
Intuitively, because the prefix tokens $x_{1}$ to $x_{N_{1}}$ do not depend on the optimized middle-stage tokens $\mathbf{x}_{N_{1}+1:N}$, the model may be unable to fully exploit the improved latent representations, potentially causing performance degradation. To test this hypothesis, we conduct an additional experiment where we optimize tokens starting from ($40\%$) into the sequence. The results, shown in Table˜13, confirm our intuition: although optimizing middle-stage tokens yields a $5.16$-point improvement over CoT on average, it still underperforms the version that optimizes the initial-stage tokens by $7.58$ points. This gap supports our hypothesis that prefix independence limits the effectiveness of optimizing middle-stage tokens.
E.8 Performance vs Fraction Ratio
Figure˜7 illustrates the performance trends of various instruction-tuned language models across different fraction ratios, which range from 0.1 to 0.8. Here, we also experimented with another model: Mistral-7B-Instruct (Jiang et al., 2023). Performance is reported as a percentage and plotted on the y-axis, while the fraction ratio is shown on the x-axis.
@@ F.1 Prompt Designation
The prompts are listed in Tables˜20, 18, 19 and 17.
@@
F.2 Backbone
We list all model backbone as follows:
F.3 Baselines
We describe all baselines in experiments as follows:
  • CoH Liu et al. (2024b): A framework which systematically aligns language models with human feedback through a structured, iterative process, enhancing their performance in complex reasoning tasks.
  • SPIN Chen et al. (2024b): A self-play fine-tuning methodology that significantly enhances the performance of weaker language models, transforming them into robust and highly capable systems competitive with stronger counterparts.
F.4 GSM8K
We provide details for GSM8K experiments as follows:
Dataset.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜21 and Table˜22, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The detailed specifications of these four evaluation prompts are provided in Tables˜17, 20, 18 and 19.The prompt structure is consistently applied across all backbones and both prompts. Notably, for Llama-3.1-8B-Instruct, Qwen2.5-1.5B-Instruct and Qwen2-7B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weights of 3, 2, and 2 assigned to format-based criteria, respectively.
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
LatentSeek (Self) Qwen2-7B-Instruct 1024 2 1 L40 0.03 Adam 0.2 bf16 10
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
Scaling.
We offer more scaling figures in Figure˜8.
(a) Qwen2.5-1.5B-Instruct
(b) Qwen2.5-7B-Instruct
(c) Qwen2.5-14B-Instruct
(d) LlaMA3.1-8B-Instruct
(e) Qwen2-7B-Instruct
F.5 MATH-500
We provide details for MATH-500 experiments as follows:
Dataset.
The MATH-500 dataset, a curated subset of the MATH benchmark, serves as a robust resource for evaluating the mathematical reasoning capabilities of machine learning models. Sourced from the HuggingFace repository at https://huggingface.co/datasets/HuggingFaceH4/MATH-500 , it comprises 500 problems derived from the original MATH benchmark developed by OpenAI (Lightman et al., 2024). Encompassing a diverse array of mathematical topics and varying difficulty levels, MATH-500 provides a comprehensive and challenging testbed for assessing model performance in mathematical problem-solving.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜23 and Table˜24, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts. Notably, for Qwen2-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using prompt 1, the weight ratios are adjusted to 1:1:1:2. Additionally, for Qwen2-7B-Instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct in LatentSeek (Self) experiment using Prompt 1, a new format reward following the methodology of DeepSeek-R1 is incorporated, with weight of 2 assigned to format-based criteria.
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
F.6 AIME2024
We provide details for AIME2024 experiments as follows:
Dataset.
Experimental Details.
For all backbones and both prompts, we use greedy decoding for inference. For the hyperparameters of LatentSeek (Self) and LatentSeek (PSRM), please refer to Table˜25 and Table˜26, respectively. The mathematical reasoning prompts we employed in the self-reward mechanism evaluate answers across four dimensions: correctness of the final answer, accuracy of problem comprehension, correctness of numerical calculations, and provision of a clear answer, weighted at 1:1:2:2, with the final score normalized to the range [-1, 0]. The prompt structure is consistently applied across all backbones and both prompts.
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
methods model max len prompt idx #GPU lr optimizer $\rho$ dtype Max Step
@@ G.1 Forward Pass FLOPs Estimation
The total FLOPs for a single forward pass is estimated by summing the operational costs of its constituent components, as detailed in Table˜27. The total FLOPs for one forward pass is calculated as follows:
@@ G.3 Total FLOPs for LatentSeek
LatentSeek operates exclusively at test time on 1319 instances. Each iteration involoves two forward passes (one for generation, one for reward evaluation) and one partial backward pass through only the LM head. From Table˜8, the average number of iterations on GSM8K is 1.27.
@@ G.4 Efficiency Threshold Analysis
Appendix H Latent Space Trajectories
Figure 9: Examples of latent-space optimization trajectories. Blue: initial latent vector; Black: intermediate steps; Green: final latent vector; Red: optimization trajectory. For more trajectories, please refer to Figure˜11 and Figure˜12
To better understand the optimization dynamics within the continuous latent space, we conduct a visualization study of latent-vector trajectories. Specifically, for each token position $t$ and optimization step $k$, we collect the corresponding latent representations $z^{(k)}_{t}$ and apply Principal Component Analysis (PCA) to project these high-dimensional vectors into a three-dimensional space.
As shown in Figure˜11 and Figure˜12 below, we visualize the projected trajectories of token latents across different positions (1st to 80th). Furthermore, Figure˜9 highlights three representative patterns of latent-space optimization. The Blue dot denotes the initial latent vector, the Black dots indicate intermediate latent states, the Green dot marks the final latent vector, and the Red line traces the optimization trajectory. As illustrated in the figures, some token-specific latents exhibit zig-zag search, others follow a consistent unidirectional trajectory, and still others trace an V-shaped search path.
Appendix I Qualitative Analysis and Case Studies
To gain a deeper understanding of model-generated responses, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021), using the Llama3.1-8B-Instruct model. These distributions are visualized in the word cloud shown in Figure˜10. The appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and unrelated proper names like “Mark” and “John”—suggests that the reasoning processes employed by language models can diverge significantly from human reasoning, despite generating superficially plausible outputs. Notably, consistent patterns emerge: the first generated tokens are often prepositions (e.g., “let”), the second tokens are primarily verbs (e.g., “find,” “solve”), and the third tokens are predominantly proper names, These findings suggest that optimal reasoning paths for language models may differ from human strategies and could be more effectively modeled in latent space.
We present several representative cases in Tables˜28, 29, 30, 31, 32, 33 and 34 to demonstrate the divergence between machine reasoning patterns and human cognition processes. These cases reveal that while our method (OURS) generates linguistically anomalous expressions such as "total downloads of downloads", "let need calculate to calculate" and "let’ll more understand it", it still achieves correct solutions. This suggests that even if the decoded tokens from our updated embeddings lack syntactic and semantic coherence, the logic and answer of the final generated output remain accurate. Notably, although the baseline model produces erroneous answers that appear fundamentally distinct from ours (OURS) in textual form, their latent representations exhibit structural proximity - through minimal adjustments to embeddings in the latent space, the original output (ORIGINAL MODEL) can be transfromed into correct solutions (OURS). These empirical evidences yield two critical implications. First, it suggests that the reasoning paths employed by machines are fundamentally distinct from those of humans. Second, while model errors may appear substantially deviant from correct answers in textual outputs, they may maintain significant proximity to valid solutions in latent space.
@@ Appendix I Qualitative Analysis and Case Studies
Figure 11: Latent-space trajectories for tokens at positions 1–40. Blue: initial latent vector; Black: intermediate steps; Green: final latent vector; Red: trajectory.
Figure 12: Latent-space trajectories for tokens at positions 41–80. Blue: initial latent vector; Black: intermediate steps; Green: final latent vector; Red: trajectory.
@@
Appendix J Computational Resources
@@
Appendix K The Use of Large Language Models (LLMs)